What is the specific heat capacity if it takes 1000 J of energy to heat 25 g of this substance by 100°C? Give your answer in J/kg°C.
1
Expert's answer
2020-04-02T10:56:01-0400

The specific heat capacity

"c=\frac{Q}{m\Delta T}=\frac{1000\:\rm J}{0.025\:\rm kg\times 100\:\rm ^{\circ}C}=400\:\rm J/(kg ^{\circ}C)."